# Anubis: Towards Reliable Cloud AI Infrastructure via Proactive Validation

### v1

@@ Abstract
We introduce Anubis, a proactive validation system for AI infrastructure that mitigates hidden degradation caused by hardware redundancies and enhances overall reliability. Anubis features a comprehensive benchmark suite, capable of evaluating individual hardware components and representing most real AI workloads. It comprises a Validator which learns benchmark criteria to clearly pinpoint defective components. Additionally, Anubis incorporates a Selector to balance validation time and issue-related penalties, enabling optimal timing for validation execution with a tailored subset of benchmarks. Through testbed evaluation and simulation, we demonstrate that Anubis can increase the mean time between incidents by up to 22.61×22.61$\times$22.61 ×. Anubis has been successfully deployed in Azure production, validating hundreds of thousands of GPUs over the last two years.
@@ 1 Introduction
To address these challenges, we introduce a proactive validation system, Anubis, featuring three key modules: (i) First, we analyze workload distributions in large-scale, multi-tenant AI clusters to identify a small number of representative workloads and parameters applicable to most customers. Based on this, we develop a comprehensive benchmark set, including end-to-end benchmarks for representative workload patterns and micro-benchmarks for individual hardware components. (ii) A Validator is designed to conduct a series of benchmarks on specified nodes. To clearly distinguish between the benchmark results of defective and functional components, the space of cumulative distribution, rather than average metrics, is used for similarity clustering to learn benchmark criteria offline. (iii) To balance the trade-off between validation time and incident coverage, benchmarks are selected based on real-time estimations of incident probability. A Selector is designed to predicts node incident through a probability model and determine when to validate on which subset of benchmarks. The selected benchmarks are executed by the Validator, generating new benchmark and defect data to periodically update criteria and probability model, allowing the entire system to evolve in tandem with the latest node statuses.
Anubis has been successfully deployed in one of the largest real-world AI infrastructure, Azure, for more than 2 years, where it identifies $10.36\%$ nodes as defects, contributing to enhanced reliability. All benchmarks in Anubis have been open-sourced on GitHub [^1] https://github.com/microsoft/superbenchmark and widely used by AI hardware vendors [1, 2] as a standard. Simulation results demonstrate that proactive validation in Anubis can increase MTBI by 22.61×22.61$\times$22.61 × and $1\times1$ compared to the absence of validation and full set validation without benchmark selection, while increasing user GPU hours by 4.81×4.81$\times$4.81 × and reducing validation time cost by $92.07\%$, respectively.
@@
  • We establish the design goals for a proactive validation system and implement Anubis to address these aims.
  • We conduct extensive evaluations of Anubis and deploy the system in a production cloud environment, demonstrating its benefits and practicality in real-world.
@@ 2.1 Massive Incidents in AI Infrastructure
Upon analyzing the related incidents, we discovered that the main reasons for high incident rate stem from three aspects: rapid-evolving AI hardware, immature software, and the cloud environment. We will explain more in the following paragraphs with real cases observed in Azure.
@@
Cloud infrastructure environments can introduce additional incidents, as they differ from vendors’ qualification environments in terms of power, temperature, and other factors. For example, tropical data centers experience more incidents due to higher temperatures. We observed a 35×35$\times$35 × increase in defective InfiniBand links with $>10^{-12}$ bit error rate in tropical data centers compared to others, leading to significantly degraded performance for training and inference. Another example is GPU throttling. Even within the same data center, different racks or locations within the same rack can exhibit varying temperatures. However, all GPUs are designed with identical cooling and heatsinks by vendors, resulting in GPUs located in warmer locations potentially experiencing thermal throttling if they cannot receive more cool air.
@@ 2.2 Observations on Hardware Redundancies
Component-wise redundancy can somewhat conceal end-to-end performance issues, but it would lead to more frequent incidents over time, diminishing the reliability of the entire infrastructure. As shown in Figure 4, we count the duration between customer-reported incidents for each node in one cluster, then calculate the mean duration between $i$th and $i+1$th incidents across all nodes that have $i+1$ incidents occurred. We find that the mean duration decreases from 719.4 hours to 151.7 hours compared between the 1st incident and the 20th incident. Furthermore, we infer the time to failure for jobs at different scales, supposing all nodes in the job have the $i$th incident occurred and share the constant failure rate.
@@
It is hard for troubleshooting to identify the root cause of each incident and restore reliability. In a distributed AI system, there are many components and tiers, some of which may not provide appropriate error information when an incident occurs. A process of elimination is required to investigate each component and tier before discovering the actual cause of failure or regression. Moreover, such workload incidents can also be closely tied to customers’ applications and may not be directly related to infrastructure issues. Additionally, due to asynchronous execution in GPUs and adaptive routing in InfiniBand networks, incidents can be non-deterministic and hard to reproduce.
@@ 3.1 Overview
Figure 7: Anubis system architecture and an example workflow.
Anubis is proposed to tackle above challenges with the following observations: (i) Despite the diversity of AI workloads, a majority can be represented by a few typical workloads, while component-wise tests can cover the rest. (ii) Instead of analyzing average metrics, all measured intermediate data (or its distribution) from a single benchmark run can be employed to establish a clear-cut boundary with maximum margins between defects or normal ones. (iii) The incident probability of a node increases gradually with use. Thus, a few benchmarks may cover most potential issues if that node passed a validation involving all benchmarks in the recent past.
@@
The above observations guide the principles applied in the Anubis system design:
@@
As depicted in Figure 7, Anubis system comprises two primary components: a Selector and a Validator. The Selector determines when and which benchmark to execute based on incoming events. It offline builds a probability model to predict nodes incident probability based on their real-time statuses, and selects an optimal benchmark subset for each validation event accordingly. The Validator conducts the benchmark execution and filters defective nodes, based on execution results and offline calculated benchmark criteria. The new node statuses and benchmark results will be continuously collected by Anubis to periodically update the offline model and criteria.
@@
The validation workflow is as follows. 1 Anubis carries out two types of validation: event-triggered and regular validation. The former is initiated by events from orchestration system, such as the node additions, node allocations for jobs. The latter is periodically checked by the Selector to validate existing nodes with high risks. 2 For each validation, Selector will first obtain statuses of related nodes, like total up time, historical incident count, MTBI of different incident types, etc., and predict incident probability with an offline trained model. 3 If the incident probability is too high for the usage, Selector will select a subset of benchmarks considering the historical benchmark coverage and validation time trade-off. 4 The subset decision will be passed to Validator. 5 Validator will then execute the selected subset of benchmarks on the corresponding nodes and get the results for each benchmark. 6 The results will be compared with Validator criteria, which is offline unsupervised learned from historical benchmark results. The node will be filtered as defective if any of its benchmark result has lower similarity with criteria than a threshold. 7 Anubis outputs the defective nodes for further handling. A repair system can be employed to maintain a defective buffer with nodes to be out for repair (OFR) and a hot buffer with repaired healthy nodes. It can swap new defective nodes with healthy ones from the hot buffer and pass to orchestration system in the runtime.
@@
The proposed Anubis validation system can be integrated into existing clusters that employ an orchestration system [62, 63, 33], requiring the following interfaces supported in the control plane:
  • Validation Events: Anubis requires specific events from the orchestration system to initiate validation, including (1) when new nodes join the cluster or cluster-wide firmware/software is upgraded, (2) when customer jobs are allocated to specific nodes, and (3) when incidents are reported by customers and particular nodes are cordoned/drained.
  • Validation Results: Anubis identifies the defective nodes from validation node set. The orchestration system can choose to use the rest of healthy nodes or find replaced nodes from a dedicated hot buffer.
@@ 3.2 Benchmark Set Design Choices
The whole validation process in Anubis relies on a comprehensive benchmark set. The benchmark set should be able to discover all known incidents and evolve when unseen incident occurs for the first time. In this section, we present how it constructs a qualified set of validation benchmarks, encompassing both end-to-end benchmarks and micro-benchmarks.
@@ 3.3 Benchmark Selection
Since the failure rate of GPU nodes changes over time due to the gradual degradation pattern, we choose to model it in neural networks without any assumptions on failure rate distribution. We apply the state-of-the-art Cox-Time model [37] for this purpose, and treat node statuses as covariates to failure rate, including time since the last incident, historical incident counts, mean time between incidents, etc., in different incident categories. The Cox-Time model is trained with historical node incident data and will estimate incident probability distribution, given node’s statuses and past incident history.
@@
Since each defective node may be covered by one or more benchmarks, this problem is NP-hard and can be considered as a variation of the 0-1 knapsack problem where different items overlap. We propose Algorithm 1 to greedily select benchmarks with $\mathcal{O}(n^{2})$ time complexity, instead of traversing all possible combinations with $\mathcal{O}(2^{n})$ complexity, based on the sub-optimal probability decrement per time unit guideline.
@@ 4 Implementation
Table 2 shows the full benchmark set chosen for Anubis. During each validation, the Selector selects a subset of benchmarks, and the Validator subsequently executes those benchmarks in two phases in sequence: the single-node phase and the multiple-node phase. In each phase, the selected benchmarks are performed in a bottom-up manner. Micro-benchmarks are conducted first to validate individual components, followed by end-to-end benchmarks to simulate customer workloads. Defective nodes are removed after each phase to ensure they do not affect subsequent benchmarks.
Networking Validation
Appendix A describes the details on the optimizations of multi-node networking validation.
Benchmark Parameter Searching
For end-to-end benchmarks, validation only needs to measure a number of iterations after performance becomes stable, rather than running into model convergence or completion. We offline search for such optimal warmup and measurement steps to reduce the validation time, Appendix B elaborates more.
Table 2: Full benchmark set in Anubis.
Micro benchmarks
End-to-end
@@
Computation
  • GPU kernel launch
  • GPU GEMM [44, 5]
  • cuBLAS kernels
  • cuDNN kernels
  • GPU burn
Communication
  • CPU latency [30]
  • GPU H2D/D2H bandwidth
  • GPU copy bandwidth
  • NVLink all-reduce
  • IB HCA loopback [55]
  • IB single-node all-reduce
Comput./Comm. Overlap
  • MatMul/all-reduce overlap
  • Sharding MatMul
Disk IO
  • FIO rand/seq read/write [10]
Multi-GPU training
  • CNN models
    • ResNet [26] 50/101/152
    • DenseNet [28] 169/201
    • VGG [59] 11/13/16/19
  • RNN models
    • LSTM
  • Transformers
    • BERT [61] base/large
    • GPT-2 [15] small/large
  • Long-running stress
    • GPT-2 large
@@
Networking
  • All pair RDMA verbs
  • GPU collective
@@
    • all-reduce
    • all-gather
    • all-to-all
Multi-node training
  • CNN models
  • RNN models
  • Transformers
  • Long-running stress
(same as above)
@@ 5.1 Experiment Settings
Anubis has been deployed in Azure production environments since Day-0 when new AI SKUs were introduced, so it is hard to perform an apple-to-apple comparison between the complete Anubis system and the original system without validation. Consequently, we divide our experiment settings into two parts: First, we gather node incident statistics and traces from on-premise clusters and use this data to simulate node incidents with and without validation, assessing the efficacy of the offline-trained incident probability model and the selection algorithm in the Selector. Second, we execute the full set of benchmarks on production clusters and two testbeds to collect benchmark results for criteria evaluation in the Validator.
@@
We deploy Anubis to the testbeds and run all benchmarks with the Validator. The environment for validation uses the exact same configurations as production clusters, including firmware versions, hypervisor settings, BIOS configurations, etc. We fixed all software versions by using a dedicated VHD image for VM and a Docker image for validation execution. The VHD image is based on Ubuntu 22.04 [11], while the Docker image is based on NVIDIA NGC container image 20.12 [48]. Both images include the necessary GPU/OFED drivers and related libraries. For the benchmark selection, we leverage PyCox [37] to build and train the probability model.

### v2

# SuperBench: Improving Cloud AI Infrastructure Reliability with Proactive Validation

@@ Abstract
We introduce SuperBench, a proactive validation system for AI infrastructure that mitigates hidden degradation caused by hardware redundancies and enhances overall reliability. SuperBench features a comprehensive benchmark suite, capable of evaluating individual hardware components and representing most real AI workloads. It comprises a Validator which learns benchmark criteria to clearly pinpoint defective components. Additionally, SuperBench incorporates a Selector to balance validation time and issue-related penalties, enabling optimal timing for validation execution with a tailored subset of benchmarks. Through testbed evaluation and simulation, we demonstrate that SuperBench can increase the mean time between incidents by up to 22.61×22.61$\times$22.61 ×. SuperBench has been successfully deployed in Azure production, validating hundreds of thousands of GPUs over the last two years.
@@ 1 Introduction
To address these challenges, we introduce a proactive validation system, SuperBench, featuring three key modules: (i) First, we analyze workload distributions in large-scale, multi-tenant AI clusters to identify a small number of representative workloads and parameters applicable to most customers. Based on this, we develop a comprehensive benchmark set, including end-to-end benchmarks for representative workload patterns and micro-benchmarks for individual hardware components. (ii) A Validator is designed to conduct a series of benchmarks on specified nodes. To clearly distinguish between the benchmark results of defective and functional components, the space of cumulative distribution, rather than average metrics, is used for similarity clustering to learn benchmark criteria offline. (iii) To balance the trade-off between validation time and incident coverage, benchmarks are selected based on real-time estimations of incident probability. A Selector is designed to predicts node incident through a probability model and determine when to validate on which subset of benchmarks. The selected benchmarks are executed by the Validator, generating new benchmark and defect data to periodically update criteria and probability model, allowing the system to evolve in tandem with the latest node statuses.
SuperBench has been successfully deployed in one of the largest real-world AI infrastructure, Azure, for more than 2 years, where it identifies $10.36\%$ nodes as defects, contributing to enhanced reliability. All benchmarks in SuperBench have been open-sourced on GitHub [^1] https://github.com/microsoft/superbenchmark and widely used by AI hardware vendors [1, 2] as a standard. Simulation results demonstrate that proactive validation in SuperBench can increase MTBI by 22.61×22.61$\times$22.61 × and $1\times1$ compared to the absence of validation and full set validation without benchmark selection, while increasing user GPU hours by 4.81×4.81$\times$4.81 × and reducing validation time cost by $92.07\%$, respectively.
@@
  • We establish the design goals for a proactive validation system and implement SuperBench to address these aims.
  • We conduct extensive evaluations of SuperBench and deploy the system in a production cloud environment, demonstrating its benefits and practicality in real-world.
@@ 2.1 Massive Incidents in AI Infrastructure
Upon analyzing the related incidents, we discovered that the main reasons for high incident rate stem from three aspects: rapid-evolving AI hardware, cloud environment, and immature software. We will explain more in the following paragraphs with real cases observed in Azure.
@@
Cloud infrastructure environments can introduce additional incidents, as they differ from vendors’ qualification environments in terms of power, temperature, and other factors. For example, data centers in tropical areas experience more incidents due to higher temperatures. We observed a 35×35$\times$35 × increase in defective InfiniBand links with $>10^{-12}$ bit error rate in data centers in tropical areas compared to data centers in higher latitudes, leading to significantly degraded performance for training and inference. Another example is GPU throttling. Even within the same data center, different racks or locations within the same rack can exhibit varying temperatures. However, all GPUs are designed with identical cooling and heatsinks by vendors, resulting in GPUs located in warmer locations potentially experiencing thermal throttling if they cannot receive more cool air.
@@ 2.2 Observations on Hardware Redundancies
Component-wise redundancy can somewhat conceal end-to-end performance issues, but it would lead to more frequent incidents over time, diminishing the reliability of the entire infrastructure. As shown in Figure 4, we count the duration between customer-reported incidents for each node in one cluster, which has 20.7k jobs per month, then calculate the mean duration between $i$th and $i+1$th incidents across all nodes that have $i+1$ incidents occurred. We find that the mean duration decreases from 719.4 hours to 151.7 hours compared between the 1st incident and the 20th incident. Furthermore, we infer the time to failure for jobs at different scales, supposing all nodes in the job have the $i$th incident occurred and share the constant failure rate.
@@
It is hard for troubleshooting to identify the root cause of each incident and restore reliability. In a distributed AI system, there are many components and tiers, some of which may not provide appropriate error information when an incident occurs. For example, some incidents suffered “uncorrectable NVLink error detected during the execution” error reported by ML framework, but the root cause turned out to be timeout in InfiniBand network rather than broken NVLink. A process of elimination is required to investigate each component and tier before discovering the actual cause of failure or regression. Moreover, such workload incidents can also be closely tied to customers’ applications and may not be directly related to infrastructure issues. Additionally, due to asynchronous execution in GPUs and adaptive routing in InfiniBand networks, incidents can be non-deterministic and hard to reproduce.
@@ 3.1 Overview
Figure 7: SuperBench system architecture and an example workflow.
SuperBench is proposed to tackle above challenges with the following observations: (i) Despite the diversity of AI workloads, a majority can be represented by a few typical workloads, while component-wise tests can cover the rest. (ii) Instead of analyzing average metrics, all measured intermediate data (or its distribution) from a single benchmark run can be employed to establish a clear-cut boundary with maximum margins between defects or normal ones. (iii) The incident probability of a node increases gradually with use. Thus, a few benchmarks may cover most potential issues if that node passed a validation involving all benchmarks in the recent past.
@@
The above observations guide the principles applied in the SuperBench system design:
@@
As depicted in Figure 7, SuperBench system comprises two primary components: a Selector and a Validator. The Selector determines when and which benchmark to execute based on incoming events. It offline builds a probability model to predict nodes incident probability based on their real-time statuses, and selects a benchmark subset based on a greedy algorithm for each validation event accordingly. The Validator conducts the benchmark execution and filters defective nodes, based on execution results and offline calculated benchmark criteria. The new node statuses and benchmark results will be continuously collected by SuperBench to periodically update the offline model and criteria.
@@
The validation workflow is as follows. 1 SuperBench carries out two types of validation: event-triggered and regular validation. The former is initiated by events from orchestration system, such as the node additions, node allocations for jobs. The latter is periodically checked by the Selector to validate existing nodes with high risks. 2 For each validation, Selector will first obtain statuses of related nodes, like total up time, historical incident count, MTBI of different incident types, etc., and predict incident probability with an offline trained model. 3 If the incident probability is too high for the usage, Selector will select a subset of benchmarks considering the historical benchmark coverage and validation time trade-off. 4 The subset decision will be passed to Validator. 5 Validator will then execute the selected subset of benchmarks on the corresponding nodes and get the results for each benchmark. 6 The results will be compared with Validator criteria, which is offline unsupervised learned from historical benchmark results. The node will be filtered as defective if any of its benchmark result has lower similarity with criteria than a threshold. 7 SuperBench outputs the defective nodes for further handling. A repair system can be employed to maintain a defective buffer with nodes to be out for repair (OFR) and a hot buffer with repaired healthy nodes. It can swap new defective nodes with healthy ones from the hot buffer and pass to orchestration system in the runtime.
@@
The proposed SuperBench validation system can be integrated into existing clusters that employ an orchestration system [63, 64, 34], requiring the following interfaces supported in the control plane:
  • Validation Events: It requires specific events from the orchestration system to initiate validation, including (1) when new nodes join the cluster or cluster-wide firmware/software is upgraded, (2) when customer jobs are allocated to specific nodes, and (3) when incidents are reported by customers and particular nodes are cordoned/drained.
  • Validation Results: SuperBench identifies the defective nodes from validation node set. The orchestration system can choose to use the rest of healthy nodes or find replaced nodes from a dedicated hot buffer.
@@ 3.2 Benchmark Set Design Choices
The whole validation process in SuperBench relies on a comprehensive benchmark set. The benchmark set should be able to discover all known incidents and evolve when unseen incident occurs for the first time. In this section, we present how it constructs a qualified set of validation benchmarks, encompassing both end-to-end benchmarks and micro-benchmarks.
@@ 3.3 Benchmark Selection
Since the failure rate of GPU nodes changes over time due to the gradual degradation pattern, we need to model it without any assumptions on failure rate distribution. The Cox-Time model [38], which is an extension of the mostly used Cox proportional hazards model [19] for time-to-event prediction in survival analysis, allows for time-varying hazard and removes the proportional assumption. We apply the state-of-the-art Cox-Time model with neural network [38] for this purpose, assume node failure rate is only related to time and some covariates, and treat node statuses as covariates to failure rate, including time since the last incident, historical incident counts, mean time between incidents, etc., in different incident categories. The Cox-Time model is trained with historical node incident data and will estimate incident probability distribution, given node’s statuses and past incident history.
@@
Since each defective node may be covered by one or more benchmarks, this problem is NP-hard and can be considered as a variation of the 0-1 knapsack problem where different items overlap. We propose Algorithm 1 to greedily select benchmarks with $\mathcal{O}(n^{2})$ time complexity, where $n$ is the number of benchmarks in benchmark set $B$, instead of traversing all possible combinations with $\mathcal{O}(2^{n})$ complexity, based on the sub-optimal probability decrement per time unit guideline.
@@ 4 Implementation
Table 2 shows the full benchmark set chosen for SuperBench. Each benchmark has pre-defined and configurable parameters, such as GEMM shapes, message sizes, model batch sizes, etc., as well as randomly generated input data. During each validation, the Selector selects a subset of benchmarks, and the Validator subsequently executes those benchmarks in two phases in sequence: the single-node phase and the multiple-node phase. In each phase, the selected benchmarks are performed in a bottom-up manner. Micro-benchmarks are conducted first to validate individual components, followed by end-to-end benchmarks to simulate customer workloads. Defective nodes are removed after each phase to ensure they do not affect subsequent benchmarks.
Table 2: Full benchmark set in SuperBench.
Micro benchmarks End-to-end
@@
Computation GPU kernel launch GPU GEMM [45, 5] cuBLAS kernels cuDNN kernels GPU burn Communication CPU latency [31] GPU H2D/D2H bandwidth GPU copy bandwidth NVLink all-reduce IB HCA loopback [56] IB single-node all-reduce Comput./Comm. Overlap MatMul/all-reduce overlap Sharding MatMul Disk IO FIO rand/seq read/write [10] Multi-GPU training CNN models ResNet [27] 50/101/152 DenseNet [29] 169/201 VGG [60] 11/13/16/19 RNN models LSTM Transformers BERT [62] base/large GPT-2 [15] small/large Long-running stress GPT-2 large
@@
Networking All pair RDMA verbs GPU collective
@@
all-reduce all-gather all-to-all
Multi-node training CNN models RNN models Transformers Long-running stress (same as above)
Networking Validation
For the distributed microbenchmark which performs pairwise scans for RDMA verbs to ensure the networking quality of InfiniBand or RoCEv2 during large-scale AI training, we propose the following two algorithms to reduce validation time:
Full Scan in $\mathcal{O}(n)$ The pairwise scans for all nodes is to ensure that the networking bandwidth between any two nodes are expected so that AI workloads with all-to-all traffic (e.g., Mixture-of-Experts models) won’t have regression. The key idea is to simultaneously run network tests for all nodes, which can avoid traffic collision by leveraging the Clos network [16] property. The problem can be formulated as follows. Given a network with ${N}$ NICs, suppose ${N}\in 2\leavevmode\nobreak\ \mathbb{N}$ and pairs are symmetric, schedule all possible $\frac{{N}\cdot({N}-1)}{2}$ NIC pairs into ${N}-1$ rounds. In each round, exactly $\frac{{N}}{2}$ pairs will run pairwise benchmarks (e.g., GPU-direct RDMA write [56], NCCL/RCCL [47, 6] 2-node all-reduce) without NIC intersection. We address this problem by leveraging the circle method in round-robin tournaments [35] as the scheduling algorithm.
Quick Scan in $\mathcal{O}(1)$ To eliminate the increasing time required for large-scale networking validation, we further propose topology-aware networking validation, which scans network links within a fixed duration, irrespective of the network scale or node number. We select node pairs such that the distance (number of hops) between each pair is 2-hop (nodes under the same ToR switch), 4-hop (nodes under the same aggregation switch), 6-hop (nodes across the core switch), and so on. For each hop number, we include all nodes without node overlap, ensuring that every node is included once and only once for the given hop. For a $k$-tier fat-tree network topology, only $k$ rounds are needed for any scale. In each round, we run the selected networking benchmarks for each hop number.
Benchmark Parameter Searching
For end-to-end benchmarks, validation only needs to measure a number of iterations after performance becomes stable, rather than running into model convergence or completion. We offline search for such optimal warmup and measurement steps to reduce the validation time. Since model structures and hyper-parameters like batch size and sequence length have already been determined based on representative workloads, the search focuses on the number of warmup steps $w$ and measurement steps $n$. It aims to minimize the total step number, i.e., benchmark running time, while maximizing the repeatability across nodes. This process is performed offline.
@@ 5.1 Experiment Settings
SuperBench has been deployed in Azure production environments since Day-0 when new AI SKUs were introduced, so it is hard to perform an apple-to-apple comparison between the complete SuperBench system and the original system without validation. Consequently, we divide our experiment settings into two parts: First, we gather node incident statistics and traces from on-premise clusters and use this data to simulate node incidents with and without validation, assessing the efficacy of the offline-trained incident probability model and the selection algorithm in the Selector. Second, we execute the full set of benchmarks on production clusters and two testbeds to collect benchmark results for criteria evaluation in the Validator.
@@
We deploy SuperBench to the testbeds and run all benchmarks with the Validator. The environment for validation uses the exact same settings as production clusters, including firmware versions, hypervisor settings, BIOS configurations, etc. We fixed all software versions by using a dedicated VHD image for VM and a Docker image for validation execution. The VHD image is based on Ubuntu 22.04 [11], while the Docker image is based on NVIDIA NGC container image 20.12 [49]. Both images include the necessary GPU/OFED drivers and related libraries. For the benchmark selection, we leverage PyCox [38] to build and train the probability model.